The CONTRAST project is over and no further updates will be made to this page. This page is retained as a record of project activities.
GV instrumentation and flight operations for CONTRAST: Schedule of daily activities and status updates
Maintenance days will start at 0800 and continue until the last team is finished or until 1600, whichever is earlier. If you have an instrument outage and need longer access hours please discuss this need with the GV project manager as soon as you become aware of such need so that RAF can schedule staff to support your activity.
When people finish maintenance on instruments, please tell the RAF crew that you are done and departing. If you leave with the intent of returning and do not tell anyone, RAF may shut dow the GV and leave if nobody else is working:.We do not want to run the aircraft power any longer than necessary as it contributes to wear on electrical systems of the aircraft, not to mention our carbon footprint.
If you need to leave for a while to get a part for example, or do something else maintenance related, please tell the RAF mechanic or technician that you will be back, and when.
Maintenance starts at 0800. Please plan to perform your work in the morning, unless it is necessary to do it later in the day, such as top off LN2 for the night, Please do not arrive at the GV at 3PM to start maintenance for a one hour task, this will not be supported. If you have an afternoon task please tell RAF PM, mechanic or technician so that we plan to support it. We may bring the power down and then return again later if the task does not require power on the GV.
Activities are ordered most recent at the top of the list.
1/29:2014: RF07, Power and access 0730, take off 1000.
1/28/2014: Maintenance day with power and access at 0800.
1/27/2014: Hard down day.
1/26/2014: Maintenance day with power and access at 0800.
1/25/2014:RF06. Power and access 0730, take off 1000. Flight into the Japanese FIR and some surface frontal area profiles.
1/24/2014: Maintenance day with power and access at 0800.
1/23/2014: Maintenance day with power and access at 0800.
1/22/2014: RF05. Power and access 0730, take off 1000. Potentially coordinated flight with BAe-146 and some convective outflow work.
1/21/2014: Maintenance day with power and access at 0800.
1/20/2014: Hard down day.
1/19/2014: RF04. Westbound survey flight. Power and access at 0730, take off at 1000. Estimated duration 7.5 hours.
1/18/2014: Maintenance day with power and access at 0800.
1/17/2014: RF03. Power and access at 0600, take off at 0900.
1/16/2014: Maintenance day with power and access at 0800.
1/15/2014: Maintenance day with power and access at 0800.
1/8/2014: Maintenance day, or TF02 if no flight on Tuesday.
1/7/2014: TF02 with t/o at around 10AM, or maintenance day.
1/6/2014: maintenance day, normal duty hours. Safety briefing repeat will take place at 3 PM in the conference room for people who missed the first one. In the PM, instrumentation status will be re-assessed to evaluate the potential for TF02 on Tuesday.
1/3/2014: Pilot proficiency flight tentatively scheduled for 0900, duration 2 hours. Short test flight planned for the afternoon, ETD 1300, duration 2 hours.
1/2/2014: Safety briefing will take place at 1500 MST in the Jeffco conference room. Cabin inspection walk-through (SAR) and FRR take place. All outstanding items have to be resolved by the end of the day on 1/2/2014. Tool control takes effect.
12/30/2013 - 1/2/2014: Instrument teams complete the last of the plumbing tasks; all instruments reach flight ready state. Please take a moment to review Safety Documents in preparation for the safety briefing. All staff going to Honolulu in support of RF01 need to make their reservation at the Hilton by the end of 12/30.
12/24/2013: Rack book requirement takes full effect. The GV has been weighed and all instrument removals or re-installations have to be recorded in the rack book.
12/23 - 12/27/2013: Power and access to the GV will be provided as needed by the RAF minimal crew. Availability of staff for supporting tasks is limited this week. Merry Christmas!
12/20/2013: Scheduled EMI test. GV will be rolled out at 0800, start of the EMI test is at 0830, expected duration 1.5 hours. After the EMI test the GV will remain on the ramp until 1230 for satcom and remote instrument control check. After that the GV will be placed in the hangar and weighed, during which time there will be no access to the cabin. This procedure is expected to take 1.5 hours, after which normal access will resume.
12/5/2013: AWAS rack is installed. Installation is on schedule. Pavel will be discussing with the instrument investigators the potential of carrying out the EMI test before the holiday season starts. This would allow to have a major milestone behind us, and plan the remaining few items (SAR, FRR, safety briefing, GV weigh) more freely.
11/26/2013: ADS header (data acquisition configuration) is installed on the GV. A few outstanding configurations will be added as the installation progresses. Calibration of the Rosemount temperature sensor will be performed. Installation is on schedule.
11/21/2013: Instrumentation teleconference. A few items from this meeting:
- Toga: has to have primary target list for species. We will need sub-list of 10-15 vars to update ADS header. Another sub-list for ground feed.
- AWAS: needs VNC installed or fixed on its DSM. Maybe add "can open" status to ADS feed.
- HARP install optics on tail on Monday, 11/25/2013
- UHSAS - at DMT, looking likely to be ready for CONTRAST; laser tuning is in progress
- Picarro, will arrive early December. This is loaner Picarro not needing the cryotrap.
- CH2O working and ready. Check UDP data feed from it.
- Br instrument. Fabricated, working on cabling. Plumbing week of 11/25. Need a red HIMIL on P11. Will not know if data collected in test flights, analyzer is in Guam. Only functional test.
- AMAX-DOAS remaining can installation.
- Remote control program - when will be available?
- Recording 4 cameras, to the ground every 1-3 min.
11/20/2013: AMAX-DOAS fiber will be installed in the GV wing today. NO-NO2-FO3 is on site to prepare / install the L3 rack.
11/19/2013: Receive AMAX-DOAS fiber, install in the left wing. Install L6 and L5 if fiber installation completed. L5 CH2O / Br rack installed.
11/18/2013: Installation of the GV infrastructure continues. L5 and L6 racks are placed in the GV but not finalized. Installation of L5 and L6 is contingent upon the delivery of the optical fiber for AMAX-DOAS, scheduled for November 19.